# Exposed extreme rays of the SONC cone

### v1

@@ 1. Introduction
In this work, we (partially) close this gap by fully characterizing the exposed extreme rays of the SONC cone.
A polynomial $f(\mathbf{x})=\sum_{{\boldsymbol{\alpha}}\in S}c_{{\boldsymbol{\alpha}}}\mathbf{x}^{{\boldsymbol{\alpha}}}+d\mathbf{x}^{{\boldsymbol{\beta}}}$, (where we use the standard shorthand for writing multivariate monomials, $\mathbf{x}^{\boldsymbol{\alpha}}\coloneqq x_{1}^{\alpha_{1}}x_{2}^{\alpha_{2}}\cdots x_{n}^{\alpha_{n}}$), where $S$ is an affinely independent subset of $2\mathbb{N}^{n}$ with ${\boldsymbol{\beta}}\in\mathbb{N}^{n}$ in the relative interior of the convex hull of $S$ and the coefficients $c_{{\boldsymbol{\alpha}}}>0$, is called a circuit polynomial; see also Section 2.1. Deciding nonnegativity of a circuit polynomial is equivalent to solving a system of linear equations, a fact that follows directly from the arithmetic-geometric inequality. The distinguishing feature of this approach is its sparsity-preserving nature, making it particularly well-suited for handling high-degree polynomials with sparse support. This motivates us to consider the SONC cone on a given finite ground set $\mathcal{A}\subseteq\mathbb{N}^{n}$ as the set of all SONC polynomials supported on $\mathcal{A}$, that is, all nonnegative combinations of circuit polynomials with “circuits” on $\mathcal{A}$ (i.e., $S\subset\mathcal{A}$ and ${\boldsymbol{\beta}}\in\mathcal{A}$). For the explicit definition, see Sections 2.1 and 2.2.
@@ 3. Unexposed rays
Proof.
@@ 4. Exposed extreme rays
Let $\mathcal{A}$ be a finite ground set $\mathcal{A}$, $f\in\mathcal{C}({\mathcal{A}})$, and suppose $l\colon\mathbb{R}[{\mathcal{A}}]\to\mathbb{R}$ is such that $r=\mathbb{R}_{+}f$ is an extreme ray of $\mathcal{C}({\mathcal{A}})$, $l(f)=0$, and for any extreme ray $q=\mathbb{R}_{+}g$ with $f\notin q$ we have $l(g)>0$.
@@
In other words, $E_{i+1}$ is constructed by removing the extreme points of $\operatorname*{conv}(E_{i})$, except for the ones that belong to $L_{0}$.
@@
It remains to prove (ii).
@@
In the case ${\boldsymbol{\alpha}}={\boldsymbol{\beta}}$ we specifically want the value $l(\mathbf{x}^{\boldsymbol{\beta}})$ to have the opposite sign to the coefficient of $f$ at $\mathbf{x}^{\boldsymbol{\beta}}$. In particular, this means that the only case when $l(\mathbf{x}^{\boldsymbol{\gamma}})$ is negative is when ${\boldsymbol{\gamma}}={\boldsymbol{\beta}}$ and $\boldsymbol{\beta}$} is odd (see Proposition 2.2, items 2 and 3). Observe that due to our choice of $\sigma$ and $\delta$, we have
@@ 5. Conclusion and open questions
A natural question is whether this approach can be extended to describe the exposed faces of SONC cones (not only rays) and whether an equally elegant combinatorial description can be achieved in that setting. In the case of the extreme rays considered in this work, it was possible to glean geometric intuition and to formulate the appropriate algebraic conjectures from studying four- and three-dimensional examples of SONC cones that can be easily visualized, however it is more difficult to use this approach for general faces, where the corresponding SONC cone may need to be higher dimensional to exhibit a sufficiently sophisticated geometric structure.
@@ Acknowledgments
The authors are grateful to the Australian Research Council for the continuing support. This work was partially supported by the ARC Discovery Project DP200100124. M.D. is supported by the ARC Discovery Early Career Award DE240100674

### v2

@@ 1. Introduction
In this work, we (partially) close this gap by fully characterizing the exposed extreme rays of the SONC cone. In contrast to the SOS cone and the cone of nonnegative polynomials, whose facial structure is amenable to the methods of algebraic geometry, the SONC cone is built as the conic hull of elementary cones, with structure governed by the combinatorics of the underlying polynomials. We use a convex geometric approach to characterize the exposed extreme rays of the SONC cone in terms of its combinatorial features.
A polynomial $f(\mathbf{x})=\sum_{{\boldsymbol{\alpha}}\in S}c_{{\boldsymbol{\alpha}}}\mathbf{x}^{{\boldsymbol{\alpha}}}+d\mathbf{x}^{{\boldsymbol{\beta}}}$, (where we use the standard shorthand for writing multivariate monomials, $\mathbf{x}^{\boldsymbol{\alpha}}\coloneqq x_{1}^{\alpha_{1}}x_{2}^{\alpha_{2}}\cdots x_{n}^{\alpha_{n}}$), where $S$ is an affinely independent subset of $2\mathbb{N}^{n}$ with ${\boldsymbol{\beta}}\in\mathbb{N}^{n}$ in the relative interior of the convex hull of $S$ and the coefficients $c_{{\boldsymbol{\alpha}}}>0$, is called a circuit polynomial; see also Section 2.1. Deciding nonnegativity of a circuit polynomial is equivalent to solving a system of linear equations, a fact that follows directly from the arithmetic-geometric inequality. The distinguishing feature of this approach is its sparsity-preserving nature, making it particularly well-suited for handling high-degree polynomials with sparse support. This motivates us to consider the SONC cone on a given finite ground set $\mathcal{A}\subseteq\mathbb{N}^{n}$ as the set of all SONC polynomials supported on $\mathcal{A}$, that is, all nonnegative combinations of circuit polynomials with “circuits” on $\mathcal{A}$ (i.e., $S\subseteq\mathcal{A}$ and ${\boldsymbol{\beta}}\in\mathcal{A}$). For the explicit definition, see Sections 2.1 and 2.2.
@@ 3. Unexposed rays
Before proving Proposition 3.1, we consider the following illustrative example based on the classical case of a SONC cone associated with the Motzkin polynomial. The affine version of the Motzkin polynomial is
$p_{M}(x,y)=1+x^{2}y^{4}+x^{4}y^{2}-3x^{2}y^{2},$
and it can be shown to be a nonnegative circuit polynomial using the SONC approach. Indeed, consider the SONC cone on the ground set that consists of the monomial degrees of $p_{M}$ (its Newton polytope is shown in Fig. 1).
[…]
Figure 1. Newton polytope of the Motzkin polynomial
There is only one possible circuit on this ground set, $(S,{\boldsymbol{\beta}})$ with $S=\{(0,0),(2,4),(4,2)\}$ and ${\boldsymbol{\beta}}=(2,2)$. Notice that for each ${\boldsymbol{\gamma}}\in S$ the conditions of Proposition 3.1 are satisfied, so we expect that the rays $\mathbb{R}_{+}$, $\mathbb{R}_{+}x^{2}y^{4}$, and $\mathbb{R}_{+}x^{4}y^{2}$ are unexposed extreme rays of the SONC cone $\mathcal{C}({\mathcal{A}})$ on the ground set $\mathcal{A}=\{(0,0),(2,2),(2,4),(4,2)\}$. Since the cone $\mathcal{C}({\mathcal{A}})$ is four-dimensional, we can visualize its compact three-dimensional slice, and witness that the corresponding extreme points are unexposed. This is illustrated in Fig. 2,
[…]
Figure 2. A three-dimensional affine slice of the four-dimensional SONC cone on the ground set of the Motzkin polynomial $p_{M}$. The coordinates $u$, $v$, and $w$ correspond to the representation of the slicing affine subspace as $L=p_{0}+\mathrm{span}\ \{p_{1},p_{2},p_{3}\}$ with $p=p_{0}+up_{1}+vp_{2}+wp_{3}$ for each $p\in L$. Here $p_{0}=2+2x^{2}y^{4}+2x^{4}y^{2}+x^{2}y^{2}$, $p_{1}=2+x^{2}y^{4}-2x^{4}y^{2}-2x^{2}y^{2}$, $p_{2}=-2+2x^{2}y^{4}+x^{4}y^{2}-2x^{2}y^{2}$, $p_{3}=1-2x^{2}y^{4}+2x^{4}y^{2}-2x^{2}y^{2}$. The constants are $c=13/2$ and $d=13/3$.
where the corresponding extreme points are visible on a slice of $\mathcal{C}({\mathcal{A}})$.
Proof of Proposition 3.1.
@@ 4. Exposed extreme rays
Let $\mathcal{A}$ be a finite ground set, $f\in\mathcal{C}({\mathcal{A}})$, and suppose $l\colon\mathbb{R}[{\mathcal{A}}]\to\mathbb{R}$ is such that $r=\mathbb{R}_{+}f$ is an extreme ray of $\mathcal{C}({\mathcal{A}})$, $l(f)=0$, and for any extreme ray $q=\mathbb{R}_{+}g$ with $f\notin q$ we have $l(g)>0$.
@@
In other words, $E_{i+1}$ is constructed by removing the extreme points of $\operatorname*{conv}(E_{i})$, except for the ones that belong to $L_{0}$ (see Fig. 3 for an illustrative example).
[…]
Figure 3. An example of a graded partition from the proof of Lemma 4.2 for $\mathcal{A}=\{{\boldsymbol{\alpha}}_{1},\dots,{\boldsymbol{\alpha}}_{13},{\boldsymbol{\beta}}\}$ and $S=\{{\boldsymbol{\alpha}}_{6},{\boldsymbol{\alpha}}_{8},{\boldsymbol{\alpha}}_{12}\}$. The sets $E_{0}$, $E_{1}$, $E_{2}$, and $E_{3}$ are shown as the outlines of the boundaries of their convex hulls. For the circuit $S^{\prime}=(\{{\boldsymbol{\alpha}}_{6},{\boldsymbol{\alpha}}_{11},{\boldsymbol{\alpha}}_{12}\},{\boldsymbol{\beta}})$ we have ${\boldsymbol{\gamma}}={\boldsymbol{\beta}}\in L_{0}$ and ${\boldsymbol{\alpha}}\in S^{\prime}\setminus S$ with ${\boldsymbol{\alpha}}={\boldsymbol{\alpha}}_{11}\in L_{2}$, while for $(\{{\boldsymbol{\alpha}}_{2},{\boldsymbol{\alpha}}_{4},{\boldsymbol{\alpha}}_{9}\},{\boldsymbol{\alpha}}_{5})$ we have ${\boldsymbol{\gamma}}={\boldsymbol{\alpha}}_{5}\in L_{2}$ and ${\boldsymbol{\alpha}}={\boldsymbol{\alpha}}_{4}\in L_{3}$.
@@
It remains to prove (ii). The subsequent case distinction is illustrated on the right-hand side of Fig. 3.
@@
In the case ${\boldsymbol{\alpha}}={\boldsymbol{\beta}}$, we specifically want the value $l(\mathbf{x}^{\boldsymbol{\beta}})$ to have the opposite sign to the coefficient of $f$ at $\mathbf{x}^{\boldsymbol{\beta}}$. In particular, this means that the only case when $l(\mathbf{x}^{\boldsymbol{\gamma}})$ is negative is when ${\boldsymbol{\gamma}}={\boldsymbol{\beta}}$ and $\boldsymbol{\beta}$} is odd (see Proposition 2.2, items 2 and 3). Observe that due to our choice of $\sigma$ and $\delta$, we have
@@ 5. Conclusion and open questions
A natural question is whether this approach can be extended to describe the exposed faces of SONC cones (not only rays) and whether a similar combinatorial description can be achieved in that setting. In the case of the extreme rays considered in this work, it was possible to glean geometric intuition and to formulate the appropriate algebraic conjectures from studying four- and three-dimensional examples of SONC cones that can be easily visualized, however it is more difficult to use this approach for general faces, where the corresponding SONC cone may need to be higher dimensional to exhibit a sufficiently sophisticated geometric structure.
@@ Acknowledgments
The authors are grateful to the Australian Research Council for the continuing support. This work was partially supported by the ARC Discovery Project DP200100124. M.D. is supported by the ARC Discovery Early Career Award DE240100674.